# Conformal Prediction for Long-Tailed Classification

### v2

@@ B.1 Optimal prediction sets for (weighted) macro-coverage
In this section, we state and prove a more formal version of Proposition˜2, of which Proposition˜1 is a special case with uniform weights.
Proposition 6 (Formal version of Proposition˜2).
@@
If thresholds $t_{\alpha}$ or $t_{\kappa}$ achieving exact equality do not exist, the optimal set remains of the form (21) but must be combined with randomization to achieve the optimal solution (that has weighted macro-coverage of exactly $1-\alpha$ or expected set size of exactly $\kappa$). See, for instance, Shao (2008, Theorem 6.1) for a formal statement of Lemma˜B.1 for this case.
@@
This result is a consequence of the Neyman-Pearson Lemma, which we state below using the formulation of Sadinle et al. (2019) (see also Casella and Berger, 2001, Theorem 8.3.12). We provide a proof of this lemma in Section˜B.2.
Lemma B.1 (Neyman and Pearson, 1933).
@@
Proof of Proposition 6(b). Let us now derive the optimal solution to the dual problem. To do so, we must rewrite the problem in a form where Lemma˜B.1 can be applied. First, observe that
@@
Applying Lemma˜B.1, combined with Remark 2, with $f(x,y)=p(x)$, $g(x,y)=\omega(y)p(x|y)$, the same measure $\mu$ as in the proof of Proposition 6(a), and $\nu=|{\mathcal{Y}}|-\kappa$ tells us that if there exists $\bar{t}_{\kappa}$ such that
@@
For completeness we give a proof of the Neyman-Pearson Lemma, as stated in Lemma˜B.1.
@@ B.3 Proof of Proposition 3
Proposition 3 (restated).
@@ B.4 Raw Fuzzy recovers Standard and Classwise CP
$=\frac{\sum_{z\in{\mathcal{Y}}}w_{\sigma}(z,y)\sum_{i=1}^{n_{z}}\mathbbm{1}\{S_{i}^{z}\leq\hat{q}\}}{\sum_{z\in{\mathcal{Y}}}n_{z}w_{\sigma}(z,y)+\omega_{\sigma}(y,y)}$
@@
The first inequality is obtained by lower bounding $\omega_{\sigma}(z,y)$ by $(1-\varepsilon)\omega_{\sigma}(y,y)$ in the numerator and upper bounding $\omega_{\sigma}(z,y)$ by $(1+\varepsilon)\omega_{\sigma}(y,y)$ in the denominator. The last inequality comes from the fact that $\hat{q}$ is the empirical quantile of the scores. By choosing $\varepsilon$ small enough, specifically,
@@ C.1 Dataset preparation
Our dataset preparation has two steps. We first identify or create train/val/test splits to replicate the standard machine learning pipeline. After obtaining train/val/test splits, we further split val into two subsets: the first for selecting the number of epochs (containing a random 30% of val) and the second for use as the calibration set ${\mathcal{D}}_{\mathrm{cal}}$ (containing the remaining 70%). This split is necessary to ensure that ${\mathcal{D}}_{\mathrm{cal}}$ is exchangeable with the test points; had we reused val for both epoch selection and calibration, this would violate exchangeability.
@@
Truncated versions. A key challenge of the long-tailed datasets we work with is that their test sets are also long-tailed, which hinders reliable class-conditional evaluation for rare classes. To address this, we create truncated versions of the datasets by removing classes with fewer than 101 examples, resulting in 330 Pl@ntNet-300K classes and 857 iNaturalist classes. For each remaining class, we assign 100 examples to test, then divide the remainder 90%/10% into train/val. For rare classes, we prioritize training set allocation, which may lead to zero-calibration examples. Specifically, to obtain Pl@ntNet-300K-truncated, we apply the above procedure to the combined train, val, and test splits of Pl@ntNet-300K. To obtain iNaturalist-2018-truncated, we apply the procedure to the combined train and val splits of the original iNaturalist dataset. [^5] For truncated datasets, we must compute marginal metrics differently when doing evaluation: due to the uniform class distribution of the test splits of these datasets, a simple average over all test examples does not reflect marginal performance on the true distribution. We estimate the marginal coverage as $\sum_{y\in{\mathcal{Y}}}\hat{p}(y)\hat{c}_{y}$ where $\hat{p}(y)$ is estimated using train and $\hat{c}_{y}$ is the empirical class-conditional coverage as defined in (16). A similar weighting procedure is used to estimate the average set size.
@@ C.2 Model training
After following the procedures described in Section˜C.1 to obtain train, val, and test splits of each dataset, we train a ResNet-50 (He et al., 2016) initialized to ImageNet pretrained weights for 20 epochs using a learning rate of 0.0001. We use train for training the neural network and the randomly selected 30% of val for computing the validation accuracy. We then select the epoch number that results in the highest validation accuracy (up to 20 epochs).
@@ Appendix D Additional Experimental Results
    We include additional methods: To ensure that we are evaluating fairly against existing procedures, we include one additional baseline method called Exact Classwise, a randomized version of classwise conformal that is designed to achieve exact (rather than at least) $1-\alpha$ coverage, as described in Appendix C.3 of Ding et al. (2023). We also include the Raw Fuzzy and Fuzzy methods we propose in Appendix A.
@@
    We evaluate on the truncated versions of Pl@ntNet-300K and iNaturalist-2018, as described in Section˜C.1.
In Section˜D.1, we provide additional results to complement Figure 3 in the main paper. In Section˜D.2, we provide results for additional methods. In Section˜D.3, we recreate Figure 3 using a ResNet-50 trained using focal loss rather than cross-entropy loss. In Section˜D.4, we do a case study of the implications of our methods for plant identification. In Section˜D.5, we visualize the score thresholds we obtain from our methods to demonstrate how they interpolate between the Standard and Classwise thresholds.
@@ D.1 Results for main methods
To aid interpretation of Figure 3, we extract the metric values for select methods for $\alpha=0.1$ and present them in Table˜3 for Pl@ntNet-300K and Table˜4 for iNaturalist-2018. We also include a comprehensive visualization of all methods on all datasets, including the truncated versions, in Figure 7.
@@
Method FracBelow50% $\downarrow$ UnderCovGap $\downarrow$ MacroCov $\uparrow$
MarginalCov
(desired $\geq$ 0.9)
Avg. set size $\downarrow$
@@
Method FracBelow50% $\downarrow$ UnderCovGap $\downarrow$ MacroCov $\uparrow$
MarginalCov
(desired $\geq$ 0.9)
Avg. set size $\downarrow$
@@ D.2 Results for additional methods
    Quantile. Recall the intuition described in the last paragraph of Section˜2, which says that we want a mapping that groups together classes with similar score distributions. To further develop this intuition, suppose that when computing $\hat{q}_{y}$, we assign non-zero weights only to classes with the same $1-\alpha$ score quantile as class $y$. Taking the $1-\alpha$ weighted quantile would then recover the $1-\alpha$ quantile of class $y$ as the number of samples with non-zero weight grows. This is because the mixture of distributions with the same $1-\alpha$ quantile has that same $1-\alpha$ quantile. Motivated by this idea, we map each class to the linearly interpolated empirical $1-\alpha$ quantile of its scores in the calibration set. This is similar to $\hat{q}_{y}^{\textsc{CW}}$ but not the same due to the linear interpolation. We chose to linearly interpolate to avoid the problem of rare classes being mapped to $\infty$, which happens if we apply no interpolation and directly apply $\mathrm{Quantile}_{1-\alpha}$ as it is defined in Section˜1.2. Linear interpolation of quantiles is described in Definition 7 of Hyndman and Fan (1996) and is the default interpolation method implemented in numpy.quantile() (Harris et al., 2020). Given a finite set $A\subset{\mathbb{R}}$ and level $\tau\in[0,1]$, let $\mathrm{LinQuantile}_{\tau}(A)$ denote the linearly interpolated $\tau$ quantile of the elements of $A$ or $s_{\max}$ if $A$ is empty, where $s_{\max}=\max_{i\in[n]}s(X_{i},Y_{i})$ is the maximum observed calibration score. The quantile projection is given by
@@ D.3 Results using focal loss
To understand how our proposed conformal prediction methods can be combined with existing strategies for dealing with long-tailed distributions, we run additional experiments where we replace the cross-entropy loss with the focal loss in our ResNet-50 training. The focal loss was proposed by Lin et al. (2017) to improve model accuracy on rare classes by modifying the cross-entropy loss. We use the PyTorch implementation of focal loss from https://github.com/AdeelH/pytorch-multi-class-focal-loss with the default parameter values of $\gamma=2$ and $\alpha=1$. The results are shown in Figure˜10 and are qualitatively similar to the cross-entropy results in Figure 3 in the main paper.
Figure 10: This figure is identical to Figure˜3 except the base model is trained using focal loss (Lin et al., 2017).
@@ D.4 Pl@ntNet case study
An additional challenge in plant identification is that visually similar species often have imbalanced labeled images, with one species significantly more represented. Standard conformal methods can suffer from occlusion, where the dominant species overshadows the rarer one at prediction time, so that the classifier always assigns low probability to the rare class. If the rare species is never included in the prediction sets, this can lead to a vicious cycle of increasing imbalance, as users simply confirm the classifier’s suggestions.
For this case study, we focus on comparing our proposed method Standard with $\mathsf{PAS}$ against Standard and Classwise when run at the $\alpha=0.1$ level. All methods have a marginal coverage guarantee, so we focus on comparing class-conditional coverage and set size between the methods. Figure˜11 shows the class-conditional coverage and average set size for the three methods. Species that are considered endangered by the International Union for Conservation of Nature (IUCN) are highlighted in red. We observe that Classwise results in huge prediction sets. On the other hand, Standard with $\mathsf{PAS}$ enhances the coverage of classes that have low coverage under Standard with $\mathsf{softmax}$ without producing huge sets. We provide visual examples of some endangered species from the Pl@ntNet-300K dataset in Figure˜12.
@@ D.5 Understanding movements in $\hat{q}_{y}$
where $s_{\mathsf{softmax}}$ denotes the $\mathsf{softmax}$ score function described in Section˜1.2. Thus, for a class $y\in{\mathcal{Y}}$, we refer to $1+\hat{q}\hat{p}(y)$ as the effective $\hat{q}_{y}$ of Standard with $\mathsf{PAS}$.

### v3

@@ B.1 Optimal prediction sets for (weighted) macro-coverage
In this section, we state and prove a more formal version of Proposition 2, of which Proposition 1 is a special case with uniform weights.
Proposition 6 (Formal version of Proposition 2).
@@
If thresholds $t_{\alpha}$ or $t_{\kappa}$ achieving exact equality do not exist, the optimal set remains of the form (21) but must be combined with randomization to achieve the optimal solution (that has weighted macro-coverage of exactly $1-\alpha$ or expected set size of exactly $\kappa$). See, for instance, Shao (2008, Theorem 6.1) for a formal statement of Lemma B.1 below for this case.
@@
This result is a consequence of the Neyman-Pearson Lemma, which we state below using the formulation of Sadinle et al. (2019) (see also Casella & Berger, 2001, Theorem 8.3.12). We provide a proof of this lemma in Section B.2.
Lemma B.1 (Neyman & Pearson, 1933).
@@
Proof of Proposition 6(b). Let us now derive the optimal solution to the dual problem. To do so, we must rewrite the problem in a form where Lemma B.1 can be applied. First, observe that
@@
Applying Lemma B.1, combined with Remark 2, with $f(x,y)=p(x)$, $g(x,y)=\omega(y)p(x|y)$, the same measure $\mu$ as in the proof of Proposition 6(a), and $\nu=|{\mathcal{Y}}|-\kappa$ tells us that if there exists $\bar{t}_{\kappa}$ such that
@@
For completeness we give a proof of the Neyman-Pearson Lemma, as stated in Lemma B.1.
@@ B.3 Proof of Proposition 3
We present Proposition 3 in two parts: the lower bound on coverage and the extreme case achieving a coverage close to this lower bound.
Proposition 3 (restated, first part).
@@
The following result provides an example of a score distribution for which a coverage of $(1-\alpha)^{2}$ is attained. For simplicity, the example is constructed as if the distribution and the quantile are exactly known, i.e., in the case of an infinite calibration set ($n=\infty$).
Proposition 3 (restated, second part).
There exists a joint distribution $\mathbb{P}$} of tuple (score, label) over $0,1]\times{\mathcal{Y}}$ with $|{\mathcal{Y}}|\geq 2$, such that for any $\tau\in(0,1)$, we have
${\mathbb{P}}\left(S\leq\tau q_{Y}^{\textsc{CW}}+(1-\tau)q\right)=(1-\alpha)^{2}\,$, (33)
where $q:=q({\mathbb{P}})$ is the (exact) $(1-\alpha)$-quantile of $S$ and for $y\in{\mathcal{Y}}$, $q_{y}^{\textsc{CW}}:=q_{y}^{\textsc{CW}}({\mathbb{P}})$ is the (exact) $(1-\alpha)$-quantile of $S|Y=y$.
Proof.
Let us define explicitly the distributions $\mathbb{P}$} for which the upper bound is achieved. The distribution of $(S,Y)\sim{\mathbb{P}}$ is defined by:
$Y\sim(1-\alpha)\delta_{y_{0}}+\alpha\delta_{y_{1}}$
$S|(Y=y_{0})\sim(1-\alpha)\delta_{0}+\alpha\delta_{1/2}\,,\quad S|(Y=y_{1})\sim\delta_{1}\,$,
for $y_{0}\neq y_{1}\in{\mathcal{Y}}$. Then $q=1/2$, $q_{y_{0}}^{\textsc{CW}}=0$ and $q_{y_{1}}^{\textsc{CW}}=1$. It follows that, for $\tau\in(0,1)$,
${\mathbb{P}}\left(S\leq\tau q_{Y}^{\textsc{CW}}+(1-\tau)q\right)$ $={\mathbb{P}}\left(S\leq\tau q_{y_{0}}^{\textsc{CW}}+(1-\tau)q|Y=y_{0}\right)(1-\alpha)$
$+{\mathbb{P}}\left(S\leq\tau q_{y_{1}}^{\textsc{CW}}+(1-\tau)q|Y=y_{1}\right)\alpha$
$={\mathbb{P}}\left(S\leq(1-\tau)/2|Y=y_{0}\right)(1-\alpha)$
$+{\mathbb{P}}\left(S\leq\tau+(1-\tau)/2|Y=y_{1}\right)\alpha.$
By plugging in the exact distributions of $S|Y=y_{0}$ and $S|Y=y_{1}$, we get
${\mathbb{P}}\left(S\leq\tau q_{Y}^{\textsc{CW}}+(1-\tau)q\right)={\mathbb{P}}\left(S=0|Y=y_{0}\right)(1-\alpha)+0=(1-\alpha)^{2}\,$,
where we have used that $(1-\tau)/2<1/2$ and $\tau+(1-\tau)/2<1$. ∎
@@ B.4 Raw Fuzzy recovers Standard and Classwise CP
$=\frac{\sum_{z\in{\mathcal{Y}}}w_{\sigma}(z,y)\sum_{i=1}^{n_{z}}\mathbbm{1}\{S_{i}^{z}\leq\hat{q}\}}{\sum_{z\in{\mathcal{Y}}}n_{z}w_{\sigma}(z,y)+w_{\sigma}(y,y)}$
@@
The first inequality is obtained by lower bounding $w_{\sigma}(z,y)$ by $(1-\varepsilon)w_{\sigma}(y,y)$ in the numerator and upper bounding $w_{\sigma}(z,y)$ by $(1+\varepsilon)w_{\sigma}(y,y)$ in the denominator. The last inequality comes from the fact that $\hat{q}$ is the empirical quantile of the scores. By choosing $\varepsilon$ small enough, specifically,
@@ C.1 Dataset preparation
Our dataset preparation has two steps. We first identify or create train/val/test splits to replicate the standard machine learning pipeline. After obtaining train/val/test splits, we further split val into two subsets: the first for selecting the number of epochs (containing a random 30% of val) and the second for use as the calibration set ${\mathcal{D}}_{\mathrm{cal}}$ (containing the remaining 70%). This split is necessary to ensure that ${\mathcal{D}}_{\mathrm{cal}}$ is exchangeable with the test points; if we reuse val for both epoch selection and calibration, this would violate exchangeability.
@@
Truncated versions. A key challenge of the long-tailed datasets we work with is that their test sets are also long-tailed, which hinders reliable class-conditional evaluation for rare classes. To address this, we create truncated versions of the datasets by removing classes with fewer than 101 examples, resulting in 330 Pl@ntNet-300K classes and 857 iNaturalist classes. For each remaining class, we assign 100 examples to test, then divide the remainder 90%/10% into train/val. For rare classes, we prioritize training set allocation, which may lead to having zero calibration examples. Specifically, to obtain Pl@ntNet-300K-truncated, we apply the above procedure to the combined train, val, and test splits of Pl@ntNet-300K. To obtain iNaturalist-2018-truncated, we apply the procedure to the combined train and val splits of the original iNaturalist dataset. [^6] For truncated datasets, we must compute marginal metrics differently when doing evaluation: due to the uniform class distribution of the test splits of these datasets, a simple average over all test examples does not reflect marginal performance on the true distribution. We estimate the marginal coverage as $\sum_{y\in{\mathcal{Y}}}\hat{p}(y)\hat{c}_{y}$ where $\hat{p}(y)$ is estimated using train and $\hat{c}_{y}$ is the empirical class-conditional coverage as defined in (16). A similar weighting procedure is used to estimate the average set size.
Classes with zero calibration examples.
Due to randomness in the data splitting, some classes end up with no calibration examples: 105 (out of 1081) in Pl@ntNet-300K, 606 (out of 8142) in iNaturalist-2018, 12 (out of 330) in truncated Pl@ntNet-300K, and 47 (out of 857) in truncated iNaturalist-2018.
@@ C.2 Model training
After following the procedures described in Section C.1 to obtain train, val, and test splits of each dataset, we train a ResNet-50 (He et al., 2016) initialized to ImageNet pretrained weights for 20 epochs using a learning rate of 0.0001. We use train for training the neural network and the randomly selected 30% of val for computing the validation accuracy. We then select the epoch number that results in the highest validation accuracy (up to 20 epochs).
@@ Appendix D Additional Experimental Results
    We include additional methods: To ensure that we are evaluating fairly against existing procedures, we include two additional baseline methods. The first is called Exact Classwise, a randomized version of classwise conformal that is designed to achieve exact (rather than at least) $1-\alpha$ coverage, as described in Appendix C.3 of Ding et al. (2023). The second is rank calibrated class-conditional conformal prediction (RC3P), proposed by Shi et al. (2024). We also include the Raw Fuzzy and Fuzzy methods we propose in Appendix A.
@@
    We evaluate on the truncated versions of Pl@ntNet-300K and iNaturalist-2018, as described in Section C.1.
In Section D.1, we provide additional results to complement Figure 3 in the main paper. In Section D.2, we provide results for additional methods. In Section D.3, we recreate Figure 3 using a ResNet-50 trained using focal loss rather than cross-entropy loss. In Section D.4, we do a case study of the implications of our methods for plant identification. In Section D.5, we visualize the score thresholds we obtain from our methods to demonstrate how they interpolate between the Standard and Classwise thresholds.
@@ D.1 Results for main methods
To aid interpretation of Figure 3, we extract the metric values for select methods for $\alpha=0.1$ and present them in Table 3 for Pl@ntNet-300K and Table 4 for iNaturalist-2018. We also include a comprehensive visualization of all methods on all datasets, including the truncated versions, in Figure 7. We remark that the two existing conformal prediction methods for many-class classification perform poorly in the long-tailed setting. The Clustered CP method from Ding et al. (2023) defaults to the score threshold from Standard CP for classes that have insufficient examples to confidently assign to a cluster, resulting in poor class-conditional coverage. Meanwhile, the RC3P method from Shi et al. (2024) suffers from a similar data splitting problem as Classwise CP, as it strongly relies on the estimation of the conditional score quantile, which can be very bad for classes with few calibration examples. Furthermore, it requires the estimation of the conditional top-k error for all classes, which is also hard when classes have few calibration examples. Moreover, as these quantities are estimated on the calibration set, the exchangeability assumption is violated which explains why the marginal coverage is sometimes below $1-\alpha$. The methods we propose strictly dominate RC3P; for a given average set size, we get much better class-conditional coverage.
@@
Method FracBelow50% $\downarrow$ UnderCovGap $\downarrow$ MacroCov $\uparrow$ MarginalCov (desired $\geq$ 0.9) Avg. set size $\downarrow$
@@
Method FracBelow50% $\downarrow$ UnderCovGap $\downarrow$ MacroCov $\uparrow$ MarginalCov (desired $\geq$ 0.9) Avg. set size $\downarrow$
@@ D.2 Results for additional methods
    Quantile. Recall the intuition described in the last paragraph of Section 2, which says that we want a mapping that groups together classes with similar score distributions. To further develop this intuition, suppose that when computing $\hat{q}_{y}$, we assign non-zero weights only to classes with the same $1-\alpha$ score quantile as class $y$. Taking the $1-\alpha$ weighted quantile would then recover the $1-\alpha$ quantile of class $y$ as the number of samples with non-zero weight grows. This is because the mixture of distributions with the same $1-\alpha$ quantile has that same $1-\alpha$ quantile. Motivated by this idea, we map each class to the linearly interpolated empirical $1-\alpha$ quantile of its scores in the calibration set. This is similar to $\hat{q}_{y}^{\textsc{CW}}$ but not the same due to the linear interpolation. We chose to linearly interpolate to avoid the problem of rare classes being mapped to $\infty$, which happens if we apply no interpolation and directly apply $\mathrm{Quantile}_{1-\alpha}$ as it is defined in Section 1.2. Linear interpolation of quantiles is described in Definition 7 of Hyndman & Fan (1996) and is the default interpolation method implemented in numpy.quantile() (Harris et al., 2020). Given a finite set $A\subset{\mathbb{R}}$ and level $\tau\in[0,1]$, let $\mathrm{LinQuantile}_{\tau}(A)$ denote the linearly interpolated $\tau$ quantile of the elements of $A$ or $s_{\max}$ if $A$ is empty, where $s_{\max}=\max_{i\in[n]}s(X_{i},Y_{i})$ is the maximum observed calibration score. The quantile projection is given by
@@ D.3 Results using focal loss
To understand how our proposed conformal prediction methods can be combined with existing strategies for dealing with long-tailed distributions, we run additional experiments where we replace the cross-entropy loss with the focal loss in our ResNet-50 training. The focal loss was proposed by Lin et al. (2017) to improve model accuracy on rare classes by modifying the cross-entropy loss. We use the PyTorch implementation of focal loss from https://github.com/AdeelH/pytorch-multi-class-focal-loss with the default parameter values of $\gamma=2$ and $\alpha=1$. The results are shown in Figure 10 and are qualitatively similar to the cross-entropy results in Figure 3 in the main paper.
Figure 10: This figure is identical to Figure 3 except the base model is trained using focal loss (Lin et al., 2017).
@@ D.4 Pl@ntNet case study
An additional challenge in plant identification is that visually similar species often have an imbalanced number of labeled images, with one species significantly more represented. Standard conformal methods can suffer from occlusion, where the dominant species overshadows the rarer one at prediction time, so that the classifier always assigns low probability to the rare class. If the rare species is never included in the prediction sets, this can lead to a vicious cycle of increasing imbalance, as users simply confirm the classifier’s suggestions.
For this case study, we focus on comparing our proposed method Standard with $\mathsf{PAS}$ against Standard and Classwise when run at the $\alpha=0.1$ level. All methods have a marginal coverage guarantee, so we focus on comparing class-conditional coverage and set size between the methods. Figure 11 shows the class-conditional coverage and average set size for the three methods. Species that are considered endangered by the International Union for Conservation of Nature (IUCN) are highlighted in red. We observe that Classwise results in huge prediction sets. On the other hand, Standard with $\mathsf{PAS}$ enhances the coverage of classes that have low coverage under Standard with $\mathsf{softmax}$ without producing huge sets. We provide visual examples of some endangered species from the Pl@ntNet-300K dataset in Figure 12.
@@ D.5 Understanding movements in $\hat{q}_{y}$
where $s_{\mathsf{softmax}}$ denotes the $\mathsf{softmax}$ score function described in Section 1.2. Thus, for a class $y\in{\mathcal{Y}}$, we refer to $1+\hat{q}\hat{p}(y)$ as the effective $\hat{q}_{y}$ of Standard with $\mathsf{PAS}$.